in 1990 the cost of a us postage stamp was 25 cents. in 2010, the cost of us postage stamp was 44 cents. in 20 years, what was the percent of increase in the cost of us postage stamps
Answer by oscargut(2103)

(44-25)/25 = 19/25 = 0.76
Answer: 76 %
If you have more questions i can see them if you contact me at: mthman@gmail.com
Thanks